If a farmer paid rs.2025 as interest over 3 years at 2% per annum to a moneylender, how much loan had the farmer taken?
1
Expert's answer
2011-11-01T12:49:31-0400
S=(1+ni)P
P=S/(1+ni)=2025/(1+0.02*3)=1910.38